# A Transferable Physics-Informed Framework for Battery Degradation Diagnosis, Knee-Onset Detection and Knee Prediction

### v1

@@ 2.5 Transfer learning-based battery degradation mode estimation
The battery degradation process is best understood and most studied in a laboratory environment, in which cycling conditions can be closely controlled and reference performance tests (RPTs) can be periodically conducted to characterize the battery degradation. With high-quality laboratory data, the hybrid physics-informed model (see Fig. 4) can be first developed in specific cycling scenarios (or source scenarios). However, different and even uncontrollable usage scenarios still pose a major challenge to deploying the trained model for online battery degradation mode estimation and phase detection in target scenarios. To address this, transfer learning (TL) is emerging as a promising strategy for transferring existing knowledge from different but related domains to a target domain in the field of advanced battery management [50].
In this study, a fine-tuning strategy is proposed to create a local model deployed in a target scenario. For the same type of battery, we assume that the parameters in Eqn. (2) remain unchanged across different usage scenarios, while the parameters in Eqn. (1) vary with battery usage scenarios. Some physical parameters in Eqn. (2) may indeed change significantly with battery aging, which is contradictory to this assumption. However, adapting the parameters in both models is computationally demanding and also increases the need of labeled data in a target scenario. Our experience with the DeepHPM also indicates that fine-tuning both the surrogate and dynamical NNs achieves performance comparable to fine-tuning the surrogate NN alone. Thus, the battery degradation modes in the target scenario could be estimated with satisfactory performance by freezing the dynamic NN $\mathcal{G}(\cdot)$ and fine-tuning the surrogate NN $\mathcal{F}(\cdot)$ using only a small amount of labeled data obtained from maintenance in the target scenario. In this way, the pre-trained DeepHPM using a large amount of data in the source scenario can be adapted to a target scenario that differs from the source.
@@ 2.6.1 Dataset description
The battery dataset used in this study was generated by Imperial College London and the Faraday Institution [40]. This high-quality open-source dataset consists of 40 lithium nickel manganese cobalt oxide (NMC 811)/graphite-$\text{SiO}_{\text{x}}$ cylindrical cells manufactured by LG Chem (model GBM50T2170, 5 Ah nominal capacity). The test aims to characterize battery degradation behaviors (i.e., capacity fade, resistance growth, and degradation mode analysis) under 15 different operating conditions (i.e., ambient temperature, SoC window, and discharge profile) throughout 5 experiments. Specifically, all the cells were charged with a 0.3C constant-current and constant-voltage (CC-CV) charging step, and then discharged with a 1C CC discharging step except for Experiment 4 in which cells were discharged with the World wide harmonized Light vehicle Test Protocol (WLTP) driving profile. Two battery health metrics, i.e., capacity (C/10 discharge, 25C) and internal resistance were measured periodically from RPTs. Moreover, three degradation modes, i.e., LLI, LAM_NE, and LAM_PE, were also estimated according to Kirkaldy et al. [40].
In this study, 6 cells in Experiment 1 and 6 cells in Experiment 5 are selected for source scenarios as they were aged under two different cycling protocols, and 6 cells in Experiment 4 are selected for the target scenario as they were aged under a dynamic cycling protocol (see Table 3). Their capacity fades are illustrated in Fig. 5. It can be seen that 6 of the cells (E1A, E1B, E1E, E1F, E4B, and E4C) have capacity knee occurrence, which is likely due to lithium plating at the negative electrode as they were cycled at low temperatures [40].
@@
Figure 5: Normalized capacity fade curves of 6 cells in Experiment 4 and 6 cells in Experiment 5 of ICL dataset. Note that capacity (C/10 discharge, 25C) from RPTs are used here.
@@ 2.6.2 Scenario-aware model development
Large amounts of battery data have been generated under well-controlled operating conditions in the past few years [51]. With these high-quality laboratory data, various battery degradation models can be developed for a range of applications throughout a battery’s life [52]. However, there is a lack of frameworks that enable the transfer of battery models to target scenarios different from those for which they were initially developed. To address this, we have proposed a scenario-aware pipeline to develop the best model using input features with predictive power generalized across different scenarios [41]. With the aid of this pipeline, the hybrid physics-informed model for battery degradation mode estimation and phase detection is first developed in the source scenarios and then deployed in the dynamic cycling target scenario, as illustrated in Fig. 6.
@@
At the offline training stage, the optimal structure of a hybrid physics-informed model is first searched for using a Bayesian hyperparameter optimization framework in the source scenarios [53]. Since there is no cell with knee occurrence in cycling protocol B but 4 cells with knee occurrence in cycling protocol A, we use 6 cells in cycling protocol A for developing the XGBoost classifier, and 6 cells in cycling protocol B for developing the DeepHPM. Specifically, the ambient temperature is first used as the criterion to classify cells into low-temperature (10C) cells, medium-temperature (25C) cells, and high-temperature (40C) cells. Then the battery data in each cycling protocol scenario is split into a training set (3 cells) and a test set (3 cells). Notably, equal ratios of low-temperature, medium-temperature, and high-temperature cells are kept in the training and test set at each split. Finally, the XGBoost classifier and DeepHPM using the feature set with the best performance over 5 train-test splits are used for developing the final hybrid physics-informed model. At the online deployment stage, there are 2 cells with knee occurrence in the dynamic cycling scenario. Therefore, whether or not a knee occurred on the capacity fade curve is used as the criterion to first classify cells into knee-occurrence cells and no-knee-occurrence cells. Then the battery data is split with 4 cells in a training set and 2 cells in a test set. It is also ensured that there is always one cell with knee occurrence and one cell without knee occurrence in the test set at each train-test split. Then different amounts of labeled data in the dynamic cycling scenario are used to fine-tune the surrogate NN $\mathcal{F}(\cdot)$ in the pre-trained hybrid physics-informed model to determine the minimum amount of data needed for satisfactory model performance in the dynamic cycling scenario. Lastly, to reduce the randomness effect, the train-test split is repeated 5 times and the averaged experimental results are reported.
@@ 2.6.3 Model performance evaluation metrics
where $\mathds{1}_{y_{i}=\hat{y}_{i}}$ is an indicator function that equals $1$ if the predicted class $\hat{y}_{i}$ is the same as the observed class $y_{i}$ of sample $i$ and 00 otherwise. TP (true positives) are the samples that have been predicted as class $k$ by the model when they actually belong to class $k$, while FP (false positives) are the samples that have been predicted as class $k$ by the model when they actually belong to other classes. Similarly, TN (true negatives) and FN (false negatives) can be defined. The precision for class $k$ measures the proportion of correctly predicted samples as class $k$ out of the total number of samples predicted as class $k$. The recall for class $k$ measures the proportion of correctly predicted samples as class $k$ out of the total number of samples that actually are in class $k$. The F1-score for class $k$ aggregates precision and recall for class $k$ into a harmonic mean of both. The harmonic mean can be used to find a trade-off between precision and recall for class $k$. The accuracy measures the proportion of correctly predicted samples of all 3 classes out of the total number of samples. Precision, recall, and F1-score are calculated for each class $k$ in the test set, while the accuracy is calculated for all 3 classes over the entire test set.
@@ 3 Results and discussion
To comprehensively evaluate the proposed transferable physic-informed framework, this section is divided into three subsections. First, the battery degradation mode estimation and phase detection performance of the proposed model are evaluated in source scenarios at the offline training stage. Secondly, battery degradation mode estimation and phase detection performance of the pre-trained model and the proposed transfer learning strategy are evaluated in the target scenario at the online deployment stage. Lastly, a case study discusses advanced battery management system functions that can be enabled in a performance digital twin.
@@ 3.1 Model evaluation in the source scenario
Table 4: Degradation mode estimation performance (RMSE) in the source scenario.
@@
Voltage-based 3-feature set 0.0080 0.0124 0.0128
Voltage-based 5-feature set 0.0066 0.0123 0.0132
Current-based 3-feature set 0.0083 0.0119 0.0132
Current-based 5-feature set 0.0079 0.0118 0.0126
Current-Voltage 17-feature set 0.0071 0.0140 0.0076
@@
Table 5: Degradation phase detection performance in the source scenario.
@@
From the RMSE results in Table 4 we conclude that the voltage-based 5-feature set (see Table 2) performs the best in LLI estimation, the current-based 5-feature set performs the best in LAM_NE estimation, and current-voltage 17-feature set performs the best in LAM_PE estimation. The time spent outside the upper and lower cut-off voltages, which correspond to overcharge or over-discharge conditions, can increase the likelihood of lithium-consuming degradation mechanisms like electrolyte decomposition, which irreversibly consumes lithium ions in the process of forming additional SEI layers as well as lithium plating. This may explain why the voltage-based 5 feature set performs the best in LLI estimation. In contrast, the current-based 5 feature set performs the best in LAM_NE estimation while the current-voltage 17 feature set performs the best in LAM_PE estimation. It can be rationalized that high currents drive rapid intercalation and deintercalation of lithium ions, which can induce alternating mechanical stress within the electrodes. Over time, these mechanical stresses can result in particle cracking and stress-driven LAM. Moreover, higher currents generate more heat inside the cell, and the resulting temperature increase can accelerate thermal-driven LAM, such as binder decomposition. Interestingly, current-based features are sufficient to achieve the best LAM estimation at the graphite-$\text{SiO}_{\text{x}}$ anode but the combination of current and voltage features is necessary to achieve the best LAM estimation at the NMC 811 cathode. However, current features alone do not provide information about the electrochemical potential of the cell and therefore cannot be independently utilized for the estimation of battery degradation modes. The voltage-based and current-voltage feature sets may be better choices for estimating degradation modes. The battery degradation phase detection performance, as measured by precision, recall, F1-score, and accuracy, are given in Table 5. The XGBoost classifier performs the best (closest to 1) in predicting Phase 1 and the least in predicting Phase 3. It can be rationalized that all the cells in the source scenario (i.e., cycling protocol A) have undergone Phase 1, and are now in either Phase 2 or 3. Consequently, the XGBoost classifier has been predominantly trained on data from Phase 1, with a smaller amount of data from Phase 2, and the least data from Phase 3. For the detection of Phase 1, the precision indicates that 99% of the time the model correctly predicted samples as Phase 1 out of the total number of samples predicted to be in Phase 1 in the test set. The recall indicates that 96% of the time the model correctly predicted samples as Phase 1 out of the total number of samples that actually are in Phase 1 in the test set. The F1-score can be interpreted as a weighted average between precision and recall for Phase 1, and the high value (97%) indicates a good trade-off between precision and recall for Phase 1. Finally, the accuracy indicates that the model correctly classified samples in the test set to each of the 3 degradation phases 96% of the time.
@@
Overall, these results indicate that the pre-trained hybrid physics-informed model is effective in estimating battery degradation modes using 5 histogram-based feature sets as well as detecting degradation phases in the source scenario. In the next subsection, the pre-trained model will be deployed in the target scenario.
@@ 3.2 Model evaluation in the target scenario
At the online deployment stage in the target scenario, the battery degradation mode estimation performance using 5 histogram-based feature sets and battery degradation phase detection performance of the pre-trained hybrid models are first evaluated on two cells in the test set, i.e., one cell with knee occurrence, and the other without knee occurrence. Then, the pre-trained models are fine-tuned using different amounts of labeled data, i.e., one cell with or without knee occurrence, or two cells with and without knee occurrence. Lastly, the fine-tuned models are evaluated using two cells in the test set. The battery degradation mode estimation results are summarized in Table 6, while the battery degradation phase detection results using estimated degradation modes as inputs are summarized in Table 7.
@@
Table 6: Degradation mode estimation performance (RMSE) in the target scenario.
@@
Feature set Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells Voltage-based, 3-feature set [0.0489,0.0433,0.0429] [0.0500,0.0477,0.0478] [0.0359,0.0325,0.0288] [0.0231,0.0266,0.0312] Voltage-based, 5-feature set [0.0479,0.0521,0.0434] [0.0516,0.0502,0.0488] [0.0361,0.0360,0.0278] [0.0203,0.0240, 0.0278] Current-based, 3-feature set [0.0531,0.0530,0.0419] [0.0494,0.0482,0.0491] [0.0422,0.0373,0.0333] [0.0351,0.0305,0.0336] Current-based, 5-feature set [0.0479,0.0461,0.0440] [0.0503,0.0485,0.0484] [0.0395,0.0320,0.0326] [0.0244,0.0263,0.0317] Current-Voltage, 17-feature set [0.0764,0.0597,0.0.0590] [0.0524,0.0510,0.0477] [0.0373,0.0406,0.0443] [0.0222,0.0220,0.0281] $[x,y,z]$ denotes RMSE values for LLI, LAM_NE, and LAM_PE, respectively. The bold values denote the minimum RMSE.
@@
Table 7: Degradation phase detection performance in the target scenario.
@@
Metric Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells Precision [0.63,0.71,0.00] [0.97,0.76,0.00] [0.96,0.83,1.00] [0.97,0.83,1.00] Recall [1.00,0.69,0.00] [0.96,0.98,0.00] [0.95,0.98,0.40] [0.96,0.98,0.39] F1-score [0.78,0.70,0.00] [0.96,0.86,0.00] [0.96,0.90,0.57] [0.96,0.90,0.57] Accuracy 67.24% 82.11% 88.19% 88.29% $[x,y,z]$ denotes each classification metric value for 3 degradation phases, respectively.
@@
According to the results in Table 6, the hybrid model developed using the voltage-based 5 feature set and then fine-tuned using data from two cells (i.e., one with and the other without knee occurrence) in the target scenario performs the best in LLI and LAM_PE estimation and the hybrid model developed using the current-voltage 17 feature set performs the best in LAM_NE estimation. To retain the correlation between current and voltage in terms of time spent, the current-voltage feature set is chosen to estimate different degradation modes in this target scenario and possibly in field applications as well. By estimating these three degradation modes using this current-voltage feature set, the battery’s degradation phase can also be monitored online. The degradation phase detection performance, as measured by precision, recall, F1-score, and accuracy, can be seen in Table 7 that the classification performance significantly improved after fine-tuning the model, i.e., accuracy improves from 67.24% to 82.11%, using one cell without knee and to 88.19% using one with knee occurrence. Using both one cell with and one cell without knee occurrence improved accuracy only slightly more, to 88.29%. Notably, the model did not successfully detect Phase 3 (i.e., precision, recall, and F1-score for Phase 3 are all 0.00) if fine-tuned using only one cell without knee occurrence but becomes capable if fined-tuned using one cell with knee occurrence.
@@
To demonstrate the effectiveness of the proposed fine-tuning strategy, we showcase the battery degradation mode estimation and phase detection results of a sample cell [E4C] before fine-tuning in Fig. 7, and after fine-tuning using one cell with knee occurrence in Fig. 8, respectively. The three degradation modes of the sample cell all begin with a square root dependence on time until they reach the knee-onset point (close to the inflection point), after which degradation modes grow exponentially. As a result, the sample cell then transfer to Phase 3. However, the pre-trained model does not predict the exponential growth of the degradation modes, and as a result, Phase 3 was not successfully detected using the estimated degradation modes (see Fig. 7). After fine-tuning the pre-trained model using one cell with knee occurrence in the target scenario, the model does not only successfully predict the exponential growth of three degradation modes, but also detect Phase 3 although with some delay. With degradation phases detected with high accuracy, the knee-onset point can also be detected online as the transition point from Phase 1 to Phase 2 (see Fig. 8). Notably, strong linear correlations were found between knee-onset and knee points identified using the curvature-based method in our previous work [37]. As illustrated in Fig. 9, we again find a strong linear correlation between knee-onset and knee ($\rho=0.962$) using this curvature-based identification method. With this strong linear correlation, online battery capacity knee prediction can be made from detected knee-onset points.
@@
Figure 9: The relationship between knee-onset and knee for 6 cells with knee occurrence.

### v2

@@ 2.5 Transfer learning-based battery degradation mode estimation
The battery degradation process is best understood and most studied in a laboratory environment, in which cycling conditions can be closely controlled and reference performance tests (RPTs) can be periodically conducted to characterize the battery degradation. With high-quality laboratory data, the hybrid physics-informed model (see Fig. 8A) can be first developed in specific cycling scenarios (or source scenarios). However, different and even uncontrollable usage scenarios still pose a major challenge to deploying the trained model for online battery degradation mode estimation and phase detection in target scenarios. To address this, transfer learning (TL) is emerging as a promising strategy for transferring existing knowledge from different but related domains to a target domain in the field of advanced battery management [51].
In this study, a fine-tuning strategy is proposed to create local models deployed in a target scenario. For the same type of battery, we assume that the parameters in Eqn. (2) remain unchanged across different usage scenarios, while the parameters in Eqn. (1) vary with battery usage scenarios. Some physical parameters in Eqn. (2) may indeed change significantly with battery aging, which is contradictory to this assumption. However, adapting the parameters in both models is computationally demanding and also increases the need of labeled data in a target scenario. Our experience with the DeepHPM also indicates that fine-tuning both the surrogate and dynamical NNs achieves performance comparable to fine-tuning the surrogate NN alone. Thus, the battery degradation modes in the target scenario could be estimated with satisfactory performance by freezing the dynamic NN $\mathcal{G}(\cdot)$ and fine-tuning the surrogate NN $\mathcal{F}(\cdot)$ using only a small amount of labeled data obtained from maintenance in the target scenario. In this way, the pre-trained DeepHPM using a large amount of data in the source scenario can be adapted to a target scenario that differs from the source.
@@ 2.6.1 Dataset description
The battery dataset was generated by Imperial College London and the Faraday Institution [41]. This high-quality open-source dataset consists of 40 lithium nickel manganese cobalt oxide (NMC 811)/graphite-$\text{SiO}_{\text{x}}$ cylindrical cells manufactured by LG Chem (model GBM50T2170, 5 Ah nominal capacity with a lower voltage cut-off of 2.5 V and an upper voltage cut-off of 4.2 V). The test aims to characterize battery degradation behaviors (i.e., capacity fade, resistance growth, and degradation mode analysis) under 15 different operating conditions (i.e., ambient temperature, SoC window, and discharge profile) throughout 5 experiments. Specifically, all the cells were charged with a 0.3C constant-current and constant-voltage (CC-CV) charging step, and then discharged with a 1C CC discharging step except for Experiment 4 in which cells were discharged with the World wide harmonized Light vehicle Test Protocol (WLTP) driving profile. Two battery health metrics, i.e., capacity (C/10 discharge, 25C) and internal resistance were measured periodically from RPTs. Moreover, three degradation modes, i.e., LLI, LAM_NE, and LAM_PE, were also quantified using pseudo-OCV measurements. Specifically, the pseudo-OCV data for degradation mode analysis was obtained using C/10 discharge-charge cycles at 25C between the voltage limits (2.5 V and 4.2 V). The C/10 discharge-charge curve of a sample cell is illustrated in Fig. 4, and the sampling interval of discharge-charge data is 10 seconds. Considering the well-controlled low-rate cycling conditions and the good quality of measured OCV data, we, therefore, took the quantified degradation modes as the ”true” degradation modes in our work. To address possible cumulative errors or amplification effects in the model output, we will incorporate uncertainty quantification into our model in future work.
Figure 4: The C/10 discharge-charge curve of a sample cell.
In this study, 6 cells in Experiment 1 and 6 cells in Experiment 5 are selected for source scenarios as they were aged under two different cycling protocols, and 6 cells in Experiment 4 are selected for the target scenario as they were aged under a dynamic cycling protocol (see Table 3). Their capacity fades are illustrated in Fig. 5. It can be seen that 6 of the cells (E1A, E1B, E1E, E1F, E4B, and E4C) have capacity knee occurrence, which is likely due to lithium plating at the negative electrode as they were cycled at low temperatures [41]. Note that the charging protocols in both source and target scenarios are the same, i.e., 0.3C CC-CV charge, while the discharging protocol in source scenarios (1C galvanostatic discharge) differs from that in the target scenario (WLTP discharge). Therefore, histogram features extracted in source scenarios cannot encode cycling conditions that lead to data leakage.
@@
Figure 5: Normalized capacity fade curves of 18 cells in the ICL dataset. Note that capacity (C/10 discharge, 25C) from RPTs are used here.
@@ 2.6.2 Scenario-aware model development
Large amounts of battery data have been generated under well-controlled operating conditions in the past few years [52]. With these high-quality laboratory data, various battery degradation models can be developed for a range of applications throughout a battery’s life [53]. However, there is a lack of frameworks that enable the transfer of battery models to target scenarios different from those for which they were initially developed. To address this, we have proposed a scenario-aware pipeline to develop the best model using input features with predictive power generalized across different scenarios [42]. With the aid of this pipeline, the hybrid physics-informed model for battery degradation mode estimation and phase detection is first developed in the source scenarios and then deployed in the dynamic cycling target scenario (see Fig. 9A).
@@
At the offline training stage, the structure of surrogate NN $\mathcal{F}(\cdot)$ is set the same as that of the dynamics NNs $\mathcal{G}(\cdot)$, and the hyperbolic tangent function is used as the activation function in the DeepHPM. Then, the optimal structure of a hybrid physics-informed model is searched for using a Bayesian hyperparameter optimization framework in the source scenarios [54]. Specifically, the DeepHPM was learned by minimizing the loss function defined in Eqn. (4). Xavier Normal was used for weight initialization, and Adam was used as the optimizer in the training process. The XGBoost classifier was learned separately by minimizing the loss function defined in Eqn. (12). Since there is no cell with knee occurrence in cycling protocol B but 4 cells with knee occurrence in cycling protocol A, we use 6 cells in cycling protocol A for developing the XGBoost classifier, and 6 cells in cycling protocol B for developing the DeepHPM. Specifically, the ambient temperature is first used as the criterion to classify cells into low-temperature (10C) cells, medium-temperature (25C) cells, and high-temperature (40C) cells. Then the battery data in each cycling protocol scenario is split into a training set (3 cells) and a test set (3 cells). Notably, equal ratios of low-temperature, medium-temperature, and high-temperature cells are kept in the training and test set at each split. Finally, the XGBoost classifier and DeepHPM using the feature set with the best performance over 5 train-test splits are used for developing the final hybrid physics-informed model. The optimal structures of the DeepHPM and XGBoost classifier can be found in Table 8 and Table 9B.
At the online deployment stage, there are 2 cells with knee occurrence in the dynamic cycling scenario. Therefore, whether or not a knee occurred on the capacity fade curve is used as the criterion to first classify cells into knee-occurrence cells and no-knee-occurrence cells. Then the battery data is split with 4 cells in a training set and 2 cells in a test set. It is also ensured that there is always one cell with knee occurrence and one cell without knee occurrence in the test set at each train-test split. Then different amounts of labeled data in the dynamic cycling scenario are used to fine-tune the surrogate NN $\mathcal{F}(\cdot)$ in the pre-trained hybrid physics-informed model to determine the minimum amount of data needed for satisfactory model performance in the dynamic cycling scenario. Lastly, to reduce the randomness effect, the train-test split is repeated 5 times and the averaged experimental results are reported.
@@ 2.6.3 Model performance evaluation metrics
where $\mathds{1}_{y_{i}=\hat{y}_{i}}$ is an indicator function that equals $1$ if the predicted class $\hat{y}_{i}$ is the same as the observed class $y_{i}$ of sample $i$ and 0 otherwise. TP (true positives) are the samples that have been predicted as class $k$ by the model when they actually belong to class $k$, while FP (false positives) are the samples that have been predicted as class $k$ by the model when they actually belong to other classes. Similarly, TN (true negatives) and FN (false negatives) can be defined. The precision for class $k$ measures the proportion of correctly predicted samples as class $k$ out of the total number of samples predicted as class $k$. The recall for class $k$ measures the proportion of correctly predicted samples as class $k$ out of the total number of samples that actually are in class $k$. The F1-score for class $k$ aggregates precision and recall for class $k$ into a harmonic mean of both. The harmonic mean can be used to find a trade-off between precision and recall for class $k$. The accuracy measures the proportion of correctly predicted samples of all three classes out of the total number of samples. Precision, recall, and F1-score are calculated for each class $k$ in the test set, while the accuracy is calculated for all three classes over the entire test set.
@@ 3 Results and discussion
To comprehensively evaluate the proposed transferable physics-informed framework, this section is divided into three subsections. First, the battery degradation mode estimation and phase detection performance of the proposed model are evaluated in source scenarios at the offline training stage. Secondly, battery degradation mode estimation and phase detection performance of the pre-trained model and the proposed transfer learning strategy are evaluated in the target scenario at the online deployment stage. Lastly, a case study discusses advanced battery management system functions that can be enabled in a performance digital twin.
@@ 3.1 Model evaluation in the source scenario
Table 4: DeepHPM degradation mode estimation performance in the source scenario.
@@
1D voltage-based 3-feature set 0.0080 0.0124 0.0128
1D voltage-based 5-feature set 0.0066 0.0123 0.0132
1D current-based 3-feature set 0.0083 0.0119 0.0132
1D current-based 5-feature set 0.0079 0.0118 0.0126
2D current-voltage 17-feature set 0.0071 0.0140 0.0076
@@
Table 5: XGBoost degradation phase detection performance in the source scenario.
@@
From the RMSE results in Table 4 we conclude that the voltage-based 5-feature set (see Table 2) performs the best in LLI estimation, the current-based 5-feature set performs the best in LAM_NE estimation, and current-voltage 17-feature set performs the best in LAM_PE estimation. The time spent outside the upper and lower cut-off voltages, which correspond to overcharge or over-discharge conditions, can increase the likelihood of lithium-consuming degradation mechanisms like electrolyte decomposition, which irreversibly consumes lithium ions in the process of forming additional SEI layers as well as lithium plating. This may explain why the voltage-based 5-feature set performs the best in LLI estimation. In contrast, the current-based 5 feature set performs the best in LAM_NE estimation while the current-voltage 17-feature set performs the best in LAM_PE estimation. It can be rationalized that high currents drive rapid intercalation and deintercalation of lithium ions, which can induce alternating mechanical stress within the electrodes. Over time, these mechanical stresses can result in particle cracking and stress-driven LAM. Moreover, higher currents generate more heat inside the cell, and the resulting temperature increase can accelerate thermal-driven LAM, such as binder decomposition. Interestingly, current-based features are sufficient to achieve the best LAM estimation for the graphite-$\text{SiO}_{\text{x}}$ anode but the combination of current and voltage features is necessary to achieve the best LAM estimation for the NMC 811 cathode. However, current features alone do not provide information about the electrochemical potential of the cell and therefore cannot be independently utilized for the estimation of battery degradation modes. The voltage-based and current-voltage feature sets may be better choices for estimating degradation modes.
The battery degradation phase detection performance, as measured by precision, recall, F1-score, and accuracy, are given in Table 5. The XGBoost classifier performs the best (closest to 1) in predicting Phase 1 and the least in predicting Phase 3. It can be rationalized that all the cells in the source scenario (i.e., cycling protocol A) have undergone Phase 1, and are now in either Phase 2 or 3. Consequently, the XGBoost classifier has been predominantly trained on data from Phase 1, with a smaller amount of data from Phase 2, and the least data from Phase 3. For the detection of Phase 1, the precision indicates that 99% of the time the model correctly predicted samples as Phase 1 out of the total number of samples predicted to be in Phase 1 in the test set. The recall indicates that 96% of the time the model correctly predicted samples as Phase 1 out of the total number of samples that actually are in Phase 1 in the test set. The F1-score can be interpreted as a weighted average between precision and recall for Phase 1, and the high value (97%) indicates a good trade-off between precision and recall for Phase 1. Finally, the accuracy indicates that the model correctly classified samples in the test set to each of the three degradation phases 96% of the time.
@@
Overall, these results indicate that the pre-trained hybrid physics-informed model is effective in estimating battery degradation modes using five histogram-based feature sets, as well as detecting degradation phases in the source scenario. In the next subsection, the pre-trained model will be deployed in the target scenario.
@@ 3.2 Model evaluation in the target scenario
At the online deployment stage in the target scenario, the battery degradation mode estimation performance using five histogram-based feature sets and battery degradation phase detection performance of the pre-trained hybrid models are first evaluated on two cells in the test set, i.e., one cell with knee occurrence, and the other without knee occurrence. Then, the pre-trained models are fine-tuned using different amounts of labeled data, i.e., one cell with or without knee occurrence, or two cells with and without knee occurrence in the target scenario. Lastly, the fine-tuned models are evaluated using two cells in the test set. The battery degradation mode estimation results are summarized in Table 6 and Table 10B, while the battery degradation phase detection results using estimated degradation modes as inputs are summarized in Table 7.
@@
Table 6: DeepHPM degradation mode estimation performance in the target scenario.
@@
Feature set Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells 1D voltage-based, 3-feature set [0.0489,0.0433,0.0429] [0.0500,0.0477,0.0478] [0.0359,0.0325,0.0288] [0.0231,0.0266,0.0312] 1D voltage-based, 5-feature set [0.0479,0.0521,0.0434] [0.0516,0.0502,0.0488] [0.0361,0.0360,0.0278] [0.0203,0.0240, 0.0278] 1D current-based, 3-feature set [0.0531,0.0530,0.0419] [0.0494,0.0482,0.0491] [0.0422,0.0373,0.0333] [0.0351,0.0305,0.0336] 1D current-based, 5-feature set [0.0479,0.0461,0.0440] [0.0503,0.0485,0.0484] [0.0395,0.0320,0.0326] [0.0244,0.0263,0.0317] 2D current-voltage, 17-feature set [0.0764,0.0597,0.0590] [0.0524,0.0510,0.0477] [0.0373,0.0406,0.0443] [0.0222,0.0220,0.0281] $[x,y,z]$ denotes RMSE values for LLI, LAM_NE, and LAM_PE, respectively. The bold values denote the minimum RMSE.
@@
Table 7: XGBoost degradation phase detection performance in the target scenario.
@@
Metric Model Pre-trained Fine-tuned with 1 cell without knee occurrence Fine-tuned with 1 cell with knee occurrence Fine-tuned with 2 cells Precision [0.63,0.71,0.00] [0.97,0.76,0.00] [0.96,0.83,1.00] [0.97,0.83,1.00] Recall [1.00,0.69,0.00] [0.96,0.98,0.00] [0.95,0.98,0.40] [0.96,0.98,0.39] F1-score [0.78,0.70,0.00] [0.96,0.86,0.00] [0.96,0.90,0.57] [0.96,0.90,0.57] Accuracy 67.24% 82.11% 88.19% 88.29% $[x,y,z]$ denotes each classification metric value for three degradation phases, respectively.
According to the results in Table 6, the hybrid model developed using the voltage-based 5 feature set and then fine-tuned using data from two cells (i.e., one with and the other without knee occurrence) in the target scenario performs the best in LLI and LAM_PE estimation and the hybrid model developed using the current-voltage 17-feature set performs the best in LAM_NE estimation. To retain the correlation between current and voltage in terms of time spent, the current-voltage 17-feature set is chosen to estimate different degradation modes in this target scenario and possibly in field applications as well. In addition, a robustness analysis is also conducted, in which we simulate two cases. In the first case, two Gaussian noises, i.e., $\mathcal{N}(0,0.04\sigma^{2}_{x})$ and $\mathcal{N}(0,0.25\sigma^{2}_{x})$, are added separately to the test data of 2D histogram 17-feature set in the target scenario. $\sigma_{x}$ denotes the standard deviation per input feature estimated in the source scenario; In the second case, considering that the 1D voltage-based or current-based 5-feature sets have been demonstrated to provide better model performance than their corresponding 1D 3-feature ones (see Table 6), the test data of four histogram features in extreme ranges (i.e., $I_{01}V_{01}$, $I_{01}V_{34}$, $I_{34}V_{01}$, and $I_{34}V_{34}$ in Table 2) in the target scenario are set to zero. The robustness analysis results are reported in Table 11B. It can be concluded from Table 11 that adding Gaussian noise makes model performance worse than missing four histogram features in extreme ranges. Thus, mitigating model performance degradation due to input noise should be prioritized in the field. Lastly, the performance of the DeepHPM degradation mode estimation is compared against that of the Gaussian process (GP) regression surrogate model. The results in the source scenario are reported in Table 12, and the results in the target scenario are reported in Table 13B. It can be concluded that the GP regression surrogate model performs better than the DeepHPM model in estimating three degradation modes in the source scenario but the GP regression surrogate model pretrained in the source scenario performs worse than the fine-tuned DeepHPM model in the target scenario. This is because the standard GP regression model lacks the mechanisms to incorporate the degradation mode dynamics defined in Eqn. (2), which makes it difficult to transfer degradation knowledge learned in the source domains to the target domain.
@@
By estimating the three degradation modes using the current-voltage 17-feature set, the battery’s degradation phase can also be monitored online. The degradation phase detection performance, as measured by precision, recall, F1-score, and accuracy, can be seen in Table 7 that the classification performance significantly improved after fine-tuning the model, i.e., accuracy improves from 67.24% to 82.11%, using one cell without knee and to 88.19% using one with knee occurrence. Using both one cell with and one cell without knee occurrence improved accuracy only slightly more, to 88.29%. Notably, the model did not successfully detect Phase 3 (i.e., precision, recall, and F1-score for Phase 3 are all 0.00) if fine-tuned using only one cell without knee occurrence but becomes capable if fined-tuned using one cell with knee occurrence.
@@
To demonstrate the effectiveness of the proposed fine-tuning strategy, we showcase the battery degradation mode estimation and phase detection results of a sample cell [E4C] before fine-tuning in Fig. 6, and after fine-tuning using one cell with knee occurrence in Fig. 7, respectively. The three degradation modes of the sample cell all begin with a square root dependence on time until they reach the knee-onset point (close to the inflection point), after which degradation modes grow exponentially. As a result, the sample cell then transfer to Phase 3. However, the pre-trained model does not predict the exponential growth of the degradation modes, and as a result, Phase 3 was not successfully detected using the estimated degradation modes (see Fig. 6). After fine-tuning the pre-trained model using one cell with knee occurrence in the target scenario, the model does not only successfully predict the exponential growth of three degradation modes, but also detect Phase 3 although with some delay. With degradation phases detected with high accuracy, the knee-onset point can also be detected online as the transition point from Phase 1 to Phase 2 (see Fig. 7). Notably, strong linear correlations were found between knee-onset and knee points identified using the curvature-based method in our previous work [38]. We again find a strong linear correlation between knee-onset and knee ($\rho=0.962$) using this curvature-based identification method. With this strong linear correlation, online battery capacity knee prediction can be made from detected knee-onset points.